an aluminum tube with a 15 inch diameter with an average speed of 150 ft/min and specific volume of


1
Expert's answer
2022-03-10T10:18:12-0500

"\dot m=\rho\dot V=\rho S\dot h=\pi\rho\frac{d^2}4 \dot h=14~\frac ts."